What is the profitability index (PI) of a project that has the following cash flows? The required return is 10.0%. Year Cash Flow 0 $ (14,000) 1 $ 3,000 2 $ 3,000 3 $ 3,000 4 $ 3,000 5 $ 3,000 6 $ 3,000 Group of answer choices 1.09 0.93 None of these are correct. 1.03 0.88
Answer #1

Time Cash Flows Present Value Factor Present Value А B = (1/1.10^n C = A*B 1 2 3 4 5 6 3,000.00 3,000.00 3,000.00 3,000.00 3,